If pKb for fluoride ion at 25 degree celsius is 10.83,the ionisation constant of hydrofluoric acid in water at this temperature is???

Grade:12

2 Answers

Soumika Hallder
6 years ago
pKb = 10.83     kb = 10^-10.83 = 1.48*10^-11
HF + H2O =  F- + H3O+
Ka * Kb = Kw
Ka = Kw/Kb = 10^-14/1.48*10^-11 = 6.76 * 10^-4
ankit singh
3 years ago
Kb = 10.83     kb = 10^-10.83 = 1.48*10^-11
HF + H2O =  F- + H3O+
Ka * Kb = Kw
Ka = Kw/Kb = 10^-14/1.48*10^-11 = 6.76 * 10^-4